OpenStudy (anonymous):

Select the equations that are parallel and perpendicular to y = 2x - 1 and that pass through the point (4, 1). parallel: y = 2x - 7 perpendicular: y = negative one over twox + 3 parallel: y = 2x + 2 perpendicular: y = negative one over twox + four and one half parallel: y = negative one over twox - 7 perpendicular: y = 2x + 3 parallel: y = -2x - 7 perpendicular: y = negative one over twox - 1 @jim_thompson5910 @Mertsj @Compassionate

OpenStudy (anonymous):

I don't remember how to do these.

OpenStudy (compassionate):

Parallel lines have the same slope. Perpendicular lines have opposite slopes. To find out what line it passes through, take (4, 1) and plug it into the each equation, if the equation is false then it doesn't run through that line. For example: If I got 3 = 4, that is false If I got: 3 = 3, then that is true

OpenStudy (anonymous):

Okay, thank you.

OpenStudy (anonymous):

Can you help me with the perpendicular line part? I have a hard time with fractions and I also have a massive headache from being online all day.

OpenStudy (compassionate):

y = 2x - 1 A perpendicular line has an opposite slope. What is the opposite of 2x?

OpenStudy (anonymous):

2y? > . <

OpenStudy (anonymous):

I know it's either the first answer or the last one.

OpenStudy (compassionate):

>2y Oh lawdy. My sides. \[2x \rightarrow -\frac{ 1 }{ 2x }\]

OpenStudy (anonymous):

I'm horrible at Algebra, I know. > . <

OpenStudy (anonymous):

Is it the first one?

OpenStudy (compassionate):

Okay, do this. The slope of a paralel line is the same Slope = 2 The slope of a perpendicular line is opposite: -2 Now! We are given the points (4, 1) Now take the equation y - y1 = m(x - x1) Then plug in the temrs y - 1 = 2(x - 4) Notice I plugged in the positive 2, which means this is the paralel line I'm finding. So, then you solve it. y - 1 = 2x - 8 y = 2x - 8 + 1 y = 2x - 7 Now, take the -2 (Perpendicular line) And do the same thing.

OpenStudy (anonymous):

Okay, so it is the first one then.

OpenStudy (compassionate):

It may be. (;

OpenStudy (anonymous):

GAH. I'm just gonna go with the first one.

OpenStudy (compassionate):

NO

OpenStudy (compassionate):

NO

OpenStudy (anonymous):

THEN THE LAST ONE.

OpenStudy (compassionate):

The paralel ne is: y = 2x - 7 Now, solve this to get your perpendicular line y - 1 = -2(x - 4)

OpenStudy (anonymous):

Would I add the one?

OpenStudy (compassionate):

Yes

OpenStudy (compassionate):

y - 1 = -2(x - 4) y - 1 = -2x + 8 y = -2x + 8 + 1

OpenStudy (anonymous):

Then divide -2x by 8?

OpenStudy (compassionate):

What the heck are ya' talking about?

OpenStudy (anonymous):

Again, I'm not so gifted in the Mathematical section. > . >

OpenStudy (compassionate):

Do you just need the answer?